# Random local access for sampling $k$-SAT solutions

### v1

@@ Abstract
We present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments to an arbitrary $k$-CNF formula $\Phi$, at exponential clause density. Our algorithm provides memory-less query access to variable assignments, such that the output variable assignments consistently emulate a single global satisfying assignment whose law is close to the uniform distribution over satisfying assignments to $\Phi$.
Such models were formally defined (for the more general task of locally sampling from exponentially sized sample spaces) in 2017 by Biswas, Rubinfeld, and Yodpinyanee, who studied the analogous problem for the uniform distribution over proper $q$-colorings. This model extends a long line of work over multiple decades that studies sublinear time algorithms for problems in theoretical computer science. Random local access and related models have been studied for a wide variety of natural Gibbs distributions and random graphical processes. Here, we establish feasiblity of random local access models for one of the most canonical such sample spaces, the set of satisfying assignments to a $k$-CNF formula.
@@ 1 Introduction
Efficiently sampling from a given exponentially sized solution space is a fundamental problem in computation. The quintessential such solution space in computer science is the collection of satisfying assignments to a given $k$-CNF (i.e., the solutions to a given SAT instance). Throughout, we let $\Phi$ denote an $n$-variable $(k,d)$-formula, which is a $k$-CNF formula such that each variable appears in at most $d$ clauses. We let $\Omega=\Omega_{\Phi}$ denote the set of satisfying assignments to $\Phi$, and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. In this work, we view $k$ as a constant and take $n\to\infty$.
For $d\leq(2ek)^{-1}2^{k}$, it has long been known (e.g., by the Lovász local lemma introduced by Erdős and Lovász [6] in 1975) that any such $k$-CNF $\Phi$ has at least one satisfying assignment. Moreover, seminal work of Moser and Tardos [23] gave a simple algorithm that computes such a satisfying assignment in expected linear time. However, a similarly basic question is far more poorly understood: when is it possible to efficiently sample an (approximately uniformly) random satisfying assignment to $\Phi$?
The first breakthrough in this direction came from Moitra [21] who gave a deterministic, fixed-parameter tractable algorithm for approximately sampling from $\mu$ provided that $d\lesssim 2^{ck}$ for $c\approx 1/60$. This was followed up by work of Feng, Guo, Yin and Zhang [8], who used a Markov chain approach to give a polynomial time algorithm for sampling from $\mu$ provided that $d\lesssim 2^{k/20}$.
A natural question about such algorithms is whether one can adapt them to more efficiently answer local questions about a random satisfying assignment in sublinear time. For variable $v\in V$, let $\mu_{v}$ denote the marginal distribution on $v$ induced by $\mu$. When $n$ is large, one might wish to sample a small handful of variables $v$ from $\mu_{v}$ in $o(n)$ time, without computing an entire $\Omega(n)$-sized satisfying assignment $\sigma\sim\mu$. Traditionally, algorithms for counting the number of satisfying assignments to a given $k$-CNF proceed by computing marginal probabilities of variable assignments, a task that can be completed given local access to a random assignment. Therefore, sublinear time algorithms for answering local questions can also yield speedups in more general counting algorithms.
Ideally, such a random local access model should provide query access to variable assignments such that the output enjoys the following properties:
@@
We give a more formal definition of random local access models in Section 2. Random local access models were formally introduced in recent work of Biswas, Rubinfeld, and Yodpinyanee [4], motivated by a long line of work over the past decades studying sublinear time algorithms for problems in theoretical computer science. The authors of [4] studied a different natural Gibbs distribution, the uniform distribution over proper $q$-colorings of a given $n$-vertex graph with maximum degree $\Delta$. By adapting previously studied classical and parallel sampling algorithms for $q$-colorings, they were able to construct a local access algorithm to a random proper $q$-coloring when $q\geq 9\Delta.$ The related problem of sampling partial information about huge random objects was pioneered much earlier in [13, 14]; further work in [24, 7] considers the implementation of different random graph models. Random local access and related (e.g., parallel computation) models have also been recently studied for several other random graphical processes and Gibbs distributions (cf. [4, 3, 22, 20]).
The above model for local access to random samples extends a rich line of work studying the more general local computation algorithms (LCAs), originally introduced by Rubinfield–Tamir–Vardi–Xie [25] and Alon–Rubinfield–Vardi–Xie [1]. A primary initial application of LCAs in [25] was locally constructing a satisfying assignment to a given $(k,d)$-formula $\Phi$ when $d\lesssim 2^{k/10}$. Similar to the desired properties of random local access, LCA implements query access to a large solution space in sublinear time using local probes. However, rather than sampling from a given distribution, LCA only provides local access to some consistent valid solution in the desired solution space that may otherwise be completely arbitrary in nature.
In this work, we present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments of an arbitrary $k$-CNF $\Phi$ at exponential clause density. The following is an informal version of our main result; a more formal version can be found at 4.7.
@@
Similar to [4], the proof of 1.1 adapts some of the algorithmic tools that came from the parallel/distributed sampling literature. The proof also builds upon the recent progress by Moitra [21] and Feng–Guo–Yin–Zhang [8] on sampling from bounded degree $k$-CNFs in polynomial time. The authors of [21, 8] both critically took advantage of a global variable marking (see 2.8) to better condition the marginal distributions of variables. Such an approach allows for a subset of the variable assignments to be sampled with ease; the resulting shattering of the solution space conditioned on such a partial assignment then allows one to efficiently complete the random satisfying assignment. These initial approaches have been extended and strengthened to nearly linear time algorithms that succeed for larger maximum variable degree in a flurry of recent work (cf. [21, 8, 18, 15, 11, 5, 10]).
@@
More recently, the work of He, Wang and Yin [16] used a recursive sampling scheme to sample $k$-CNF solutions. Their work immediately gives a sublinear (in fact, expected constant time) algorithm for sampling the assignment of a single variable one time; however, this work does not immediately extend out of the box to a random local access model that enjoys the desired consistency and memory-less properties. Recursive sampling schemes have also been recently used to make progress on analyzing and designing fast sampling algorithms for a variety of Gibbs distributions (cf. [16, 5, 2]).
@@
An immediate roadblock to adapting global sampling strategies to random local access to $k$-CNF solutions is that the vast majority of the aforementioned existing algorithms critically use some global information – like a variable marking, particular ordering of variables, or other global state compression – as an input to make algorithmic decisions, necessitating a substantive departure from these approaches for any hope of local access adaptation. In this work, we overcome these obstacles by carefully selecting and sampling a small number of auxiliary variable assignments on an as-needed basis, and showing that bad cases can be reduced to small brute force calculations after a sublinear number of steps. Thus our proof of 1.1 requires a novel adaptation of sampling techniques that require global information or complete passes of the variable set to the local setting, showing that we can perform analogous operations locally or circumvent them altogether in the course of locally sampling a variable assignment.
@@ Acknowledgements
We would like to thank Kuikui Liu and Ronitt Rubinfeld for helpful discussions, suggestions, and feedback on an early draft. NM was supported by the Hertz Graduate Fellowship and the NSF Graduate Research Fellowship Program #2141064.
@@ 2 Preliminaries
Throughout, we let $\Phi=(V,\mathcal{C})$ be a $k$-CNF on variable set $V=\{v_{1},\ldots,v_{n}\}$ with associated collection of clauses $\mathcal{C}=\{C_{1},\ldots,C_{m}\}$. We further assume that $\Phi$ is a $(k,d)$-formula; that is, each variable $v_{i}$ appears in at most $d$ clauses. For every clause $C\in\mathcal{C}$, let $\operatorname{vbl}(C)$ denote the collection of variables in the clause $C$. We further define $\Delta:=\max_{C\in\mathcal{C}}|\{C^{\prime}\in\mathcal{C}:\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset\}|$, so in particular $\Delta\leq kd$.
In the regime we work in, we assume $k$ is a large fixed constant and view $n\to\infty.$ We use $f\lesssim g$ to denote that there is some constant $C$ (not depending on $k$) such that $f\leq Cg$. We also use the standard asymptotic notation ($O$, $o$, $\Omega$, $\omega$, $\Theta$), where when not specified, we assume these are in the $n\to\infty$ limit and in these notations, we ignore any dependencies on $k$. We use $\mathsf{Law}(X)$ to denote the underlying distribution of a random variable $X$.
We let $\Omega=\Omega_{\Phi}\subseteq\{0,1\}^{n}$ denote the collection of satisfying assignments to $\Phi$ and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. Throughout, we suppress the $\Phi$ subscripts when the formula is clear from context. We now introduce a few more concepts associated to a $k$-CNF formula that will be used later.
Definition 2.1 (Dependency graph).
Given a $(k,d)$-formula $\Phi$, we define the following auxiliary (hyper)graphs.
  1. 1.
    Let $H_{\Phi}=(V,\mathcal{E})$ be the dependency hypergraph (with multiple edges allowed), where $V$ is the set of variables and =$\mathcal{E}=\{\operatorname{vbl}(C):C\in\mathcal{C}(\Phi)\}$ is the collection of clauses of $\Phi$ viewed as $k$-sets of variables;
  2. 2.
    $G_{\Phi}$ be the clause dependency graph of $\Phi$, where $V(G_{\Phi}):=\mathcal{C}$ is the set of clauses of $\Phi$, and two clauses $C,C^{\prime}\in\mathcal{C}$ are adjacent in $G_{\Phi}$ if they share a variable.
@@
$\mathcal{Q}^{*}:=\{0,1,\star,\circ\}^{V}$
denote the space of partial assignments with the following symbology. Given a partial assignment $\sigma\in\mathcal{Q}^{*}$, each variable $v\in V$ is classified as follows:
@@
  • $\sigma(v)=\star$ means that $v$ is just accessed by the algorithm but unassigned yet with any value;
  • $\sigma(v)=\circ$ means that $v$ is unaccessed by the algorithm and hence unassigned with any value.
We sometimes use $\emptyset$ to denote the empty assignment (i.e., $\sigma(v)=\circ$ for all $v\in V$). For every $\sigma\in\mathcal{Q}^{*}$, we use $\Lambda(\sigma)\subseteq V$ to denote the set of variables that are assigned and accessed in $\sigma$, namely
$\Lambda(\sigma):=\{v\in V:\sigma(v)\in\{0,1\}\}.$
Given $\sigma\in\mathcal{Q}^{*}$, variable $v\in V$ and $x\in\{0,1,\star,\circ\}$, we further denote by $\sigma_{v\leftarrow x}$ the partial assignment obtained from $\sigma$ by replacing $\sigma(v)$ with $x$. We say that $\sigma$ is feasible if it can be extended to a satisfying assignment to $\Phi$.
@@
Let $\Phi$ be a $(k,d)$-formula. Given a partial assignment $\sigma$ on $\Lambda:=\Lambda(\sigma)$, let $\Phi^{\sigma}$ be the result of simplifying $\Phi$ under $\sigma$. That is, define $\Phi^{\sigma}:=(V^{\sigma},\mathcal{C}^{\sigma})$ where
  • $V^{\sigma}=V\setminus\Lambda$,
  • $\mathcal{C}^{\sigma}$ is obtained from $\mathcal{C}$ by removing all clauses that have been satisfied under $\sigma$ and removing any appearance of variables in $\Lambda$.
Let $H_{\Phi}^{\sigma}$ be the associated (not necessarily $k$-uniform) hypergraph to $\Phi^{\sigma}$, and let $G_{\Phi}^{\sigma}$ be the associated clause dependency graph to $\Phi^{\sigma}$. For variable $v\in V\setminus\Lambda$, let $\Phi_{v}^{\sigma}$ denote the connected component of $\Phi^{\sigma}$ to which $v$ belongs.
@@
When $S=\{v\}$ is a single vertex, we let $\mu_{v}=\mu_{\{v\}}$. Furthermore, given some partial assignment $\sigma\in\{0,1\}^{\Lambda}$ for $\Lambda\subseteq V$, if $S\cap\Lambda=\emptyset$, we let $\mu_{S}^{\sigma}(\cdot):=\mu_{S}(\cdot\mid\sigma)$ be the marginal distribution on $S$ conditioned on the partial assignment $\sigma$. When $S=V\setminus\Lambda$, we simply use $\mu^{\sigma}$ to denote $\mu^{\sigma}_{V\setminus\Lambda}$.

### v2

@@ Abstract
We present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments to an arbitrary $k$-SAT formula $\Phi$, at exponential clause density. Our algorithm provides memory-less query access to variable assignments, such that the output variable assignments consistently emulate a single global satisfying assignment whose law is close to the uniform distribution over satisfying assignments to $\Phi$.
Such models were formally defined (for the more general task of locally sampling from exponentially sized sample spaces) in 2017 by Biswas, Rubinfeld, and Yodpinyanee, who studied the analogous problem for the uniform distribution over proper $q$-colorings. This model extends a long line of work over multiple decades that studies sublinear time algorithms for problems in theoretical computer science. Random local access and related models have been studied for a wide variety of natural Gibbs distributions and random graphical processes. Here, we establish feasibility of random local access models for one of the most canonical such sample spaces, the set of satisfying assignments to a $k$-SAT formula.
@@ 1 Introduction
Efficiently sampling from an exponentially sized solution space is a fundamental problem in computation. The quintessential such problem is sampling a uniformly random satisfying assignment to a given $k$-SAT formula, the setting of this work. Throughout, we let $\Phi$ denote an $n$-variable $(k,d)$-formula, which is a $k$-SAT formula in conjunctive normal form (which we also call a $k$-CNF) such that each variable appears in at most $d$ clauses. We let $\Omega=\Omega_{\Phi}$ denote the set of satisfying assignments to $\Phi$, and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. We consider the regime where $k$ is a (large) constant and take $n\to\infty$.
For $d\leq(2ek)^{-1}2^{k}$, it has long been known (e.g., by the Lovász local lemma introduced in [7]) that any such $k$-SAT formula $\Phi$ has at least one satisfying assignment. Moreover, seminal work of Moser and Tardos [26] gave a simple algorithm that computes such a satisfying assignment in expected linear time. However, a similarly basic question in the sampling regime is far more poorly understood: when is it possible to efficiently sample an (approximately uniformly) random satisfying assignment to $\Phi$?
The first breakthrough in this direction came from Moitra [24] who gave a deterministic, fixed-parameter tractable algorithm for approximately sampling from $\mu$ provided that $d\lesssim 2^{ck}$ for $c\approx 1/60$. This was followed up by work of Feng, Guo, Yin and Zhang [10], who used a Markov chain approach to give a polynomial time algorithm for sampling from $\mu$ provided that $d\lesssim 2^{k/20}$. Recent works [18, 29] have given state of the art results for sampling from $k$-CNFs, when $d\lessapprox 2^{k/4.82}$ (in fact for large domain sizes, [29] gives essentially tight bounds on sampling from atomic constraint satisfaction problems). There has also been a tremendous amount of work sampling from random $k$-CNFs, recently essentially resolved by work of [5], building on previous progress of [13, 6, 12, 19].
A natural question about such algorithms to sample from $k$-CNFs is whether one can adapt them to more efficiently answer local questions about a random satisfying assignment in sublinear time. For variable $v$, let $\mu_{v}$ denote the marginal distribution on $v$ induced by $\mu$. When $n$ is large, one might wish to sample a small handful of variables $S$ from their marginal distributions $\mu_{S}$ in $o(n)$ time, without computing an entire $\Omega(n)$-sized satisfying assignment $\sigma\sim\mu$. Further, many traditional algorithms for counting the number of satisfying assignments to a given $k$-SAT formula proceed by computing marginal probabilities of variable assignments, a task that can be completed given local access to a random assignment. Therefore, sublinear time algorithms for answering local questions can also yield speedups in more general counting algorithms.
Ideally, a random local access model should provide query access to variable assignments such that the output enjoys the following properties:
@@
We give a more precise definition of random local access models in Section 2. Random local access models were formally introduced in work of Biswas, Rubinfeld, and Yodpinyanee [4], motivated by a long line of work over the past decades studying sublinear time algorithms for problems in theoretical computer science. The authors of [4] studied a different natural Gibbs distribution, the uniform distribution over proper $q$-colorings of a given $n$-vertex graph with maximum degree $\Delta$. By adapting previously studied classical and parallel sampling algorithms for $q$-colorings, they were able to construct a local access algorithm to a random proper $q$-coloring when $q\geq 9\Delta.$ The related problem of sampling partial information about huge random objects was pioneered much earlier in [15, 16]; further work in [27, 8] considers the implementation of different random graph models. Random local access and related (e.g., parallel computation) models have also been recently studied for several other random graphical processes and Gibbs distributions (cf. [4, 3, 25, 23]).
The above model for local access to random samples extends a rich line of work studying local computation algorithms (LCAs), originally introduced in works of Rubinfield, Tamir, Vardi, and Xie [28] and Alon, Rubinfield, Vardi, and Xie [1]. A primary initial application of LCAs in [28] was locally constructing a satisfying assignment to a given $(k,d)$-formula $\Phi$ when $d\lesssim 2^{k/10}$. Similar to the desired properties of random local access, LCA implements query access to a large solution space in sublinear time using local probes. However, rather than sampling from a given distribution, LCA only provides local access to some consistent valid solution in the desired solution space that may otherwise be completely arbitrary in nature.
In this work, we present a polylogarithmic time algorithm that gives random local access to the uniform distribution over satisfying assignments of an arbitrary $k$-SAT formula $\Phi$ at exponential clause density. The following is an informal version of our main result; a more formal version can be found at 5.3.
@@
Similar to [4], the proof of 1.1 adapts some of the algorithmic tools used to study parallel and distributed sampling. The proof also builds upon the work of Moitra [24] and Feng–Guo–Yin–Zhang [10] on sampling from bounded degree $k$-CNFs in polynomial time. The authors of [24, 10] both critically took advantage of a global variable marking (see 2.9) to better condition the marginal distributions of variables. Such an approach allows for a subset of the variable assignments to be sampled with ease; the resulting shattering of the solution space conditioned on such a partial assignment then allows one to efficiently complete the random satisfying assignment. These initial approaches have been extended and strengthened to nearly linear time algorithms that succeed for larger maximum variable degree in a flurry of recent work (c.f. [24, 10, 20, 17, 13, 6, 12, 29, 5]).
@@
Recently, Feng, Guo, Wang, Wang, and Yin [9] used a recursive sampling scheme to simulate the systematic scan projected Glauber dynamics via a strategy termed coupling towards the past, which they used to derandomise several MCMC algorithms for sampling from CSPs. Additionally, recent work of He, Wang and Yin [18] used a recursive sampling scheme to sample $k$-SAT solutions. Their work immediately gives a sublinear (in fact, expected constant time) algorithm for sampling the assignment of a single variable one time; however, this work does not immediately extend out of the box to a random local access model that enjoys the desired consistency and memory-less properties if multiple variables are sampled. Recursive sampling schemes have also been recently used to make progress on analyzing and designing fast sampling algorithms for a variety of Gibbs distributions (cf. [18, 6, 2]). As noted earlier, such schemes have been particularly useful for derandomising and constructing parallel and distributed versions of many popular MCMC algorithms for sampling solutions to CSPs [9, 11, 17, 18].
@@
An immediate roadblock to adapting such global or parallel sampling strategies to random local access to $k$-SAT solutions is that the vast majority of the aforementioned existing algorithms critically either use some global information – like a variable marking, particular ordering of variables, or other global state compression – as an input to make algorithmic decisions or postpone sampling of certain problematic variable assignments until a linear number of other choices are made. Both these issues necessitate a substantive departure from these approaches for any hope of local access adaptation. In this work, we overcome these obstacles by adapting the coupling towards the past strategy used in [9] to derandomise MCMC in conjunction with a local implementation of the variable marking strategy introduced in [24].
We use these algorithms to carefully select and sample a small number of auxiliary variable assignments on an as-needed basis, showing that bad cases can be reduced to small calculations after a sublinear number of steps. Our proof of 1.1 requires a novel adaptation of sampling techniques to avoid requiring global information or complete passes of the variable set; we show that we can perform analogous operations locally or leverage local uniformity properties to circumvent them altogether in the course of locally sampling a variable assignment. Importantly, we demonstrate how to execute local sampling in an oblivious, consistent fashion, so that the algorithm need not retain any memory and that variables need not be sampled in any particular order.
@@ Acknowledgements
We would like to thank Kuikui Liu and Ronitt Rubinfeld for helpful discussions and feedback on a draft. We would like to thank the anonymous referee for identifying an issue with a previous version of this article. NM was supported by the Hertz Graduate Fellowship and the NSF Graduate Research Fellowship Program #2141064. This material includes work supported by the NSF Grant DMS-1928930, while the authors were in residence at SLMath during the Spring 2025 semester.
@@ 2 Preliminaries
Throughout, let $\Phi=(V,\mathcal{C})$ be a $k$-CNF on variable set $V=\{v_{1},\ldots,v_{n}\}$ with associated collection of clauses $\mathcal{C}$. We further assume that $\Phi$ is a $(k,d)$-formula; that is, each variable $v_{i}$ appears in at most $d$ clauses. For every clause $C\in\mathcal{C}$, let $\operatorname{vbl}(C)$ denote the collection of variables in the clause $C$. We further define $\Delta:=\max_{C\in\mathcal{C}}|\{C^{\prime}\in\mathcal{C}:\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset\}|$, so in particular $\Delta\leq kd$.
In the regime we work in, we assume $k$ is a large fixed constant and view $n\to\infty.$ We use $f\lesssim g$ to denote that there is some constant $C$ (not depending on $k$) such that $f\leq Cg$. We also use the standard asymptotic notation ($O$, $o$, $\Omega$, $\omega$, $\Theta$), where when not specified, we assume these are in the $n\to\infty$ limit. We use $\mathsf{Law}(X)$ to denote the underlying distribution of a random variable $X$.
We let $\Omega=\Omega_{\Phi}\subseteq\{0,1\}^{n}$ denote the set of satisfying assignments to $\Phi$ and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. We suppress the $\Phi$ subscripts when the formula is clear from context. We introduce a few more concepts associated to a $k$-SAT instance that will be used later.
Definition 2.1.
Given probability distributions $\nu_{1},\nu_{2}$ over $\Omega$, the total variation distance is
$d_{\operatorname{\mathsf{TV}}}(\nu_{1},\nu_{2}):=\frac{1}{2}\sum_{\omega\in\Omega}|\nu_{1}(\omega)-\nu_{2}(\omega)|.$
If we have a random variable $X$ with $\textsf{Law}(X)=\nu$, we may write $d_{\operatorname{\mathsf{TV}}}(\mu,X)$ instead of $d_{\operatorname{\mathsf{TV}}}(\mu,\nu)$ in a slight abuse of notation.
Definition 2.2 (Dependency hypergraph).
Given a $(k,d)$-formula $\Phi$, let $H_{\Phi}=(V,\mathcal{E})$ be the dependency hypergraph (with multiple edges allowed), where $V$ is the set of variables and =$\mathcal{E}=\{\operatorname{vbl}(C):C\in\mathcal{C}\}$ is the collection of clauses of $\Phi$ viewed as $k$-sets of variables.
@@
$\mathcal{Q}^{*}:=\bigcup_{\Lambda\subseteq V}\{0,1,\perp\}^{\Lambda}$
denote the space of partial assignments with the following symbology. Given a partial assignment $\sigma\in\{0,1,\perp\}^{\Lambda}$ on some $\Lambda\subseteq V$, each variable $v\in\Lambda$ is classified as follows:
@@
  • $\sigma(v)=\perp$ means that $v$ is accessed by the algorithm but unassigned yet with any value.
We sometimes use $\emptyset$ to denote the empty assignment (i.e., $\Lambda$ is the empty set). We say that $\sigma$ is feasible if it can be extended to a satisfying assignment to $\Phi$.
@@
Let $\Phi$ be a $(k,d)$-formula. Given a partial assignment $\sigma$ on $\Lambda\subseteq V$, let $\Phi^{\sigma}$ be the result of simplifying $\Phi$ under $\sigma$. That is, define $\Phi^{\sigma}:=(V^{\sigma},\mathcal{C}^{\sigma})$ where
  • $V^{\sigma}=V\setminus\{v\in\Lambda:\sigma(v)\in\{0,1\}\}$,
  • $\mathcal{C}^{\sigma}$ is obtained from $\mathcal{C}$ by removing all clauses that have been satisfied under $\sigma$, and removing any appearance of variables that are assigned 0 or 1 by $\sigma$.
Let $H_{\Phi}^{\sigma}$ be the associated (not necessarily $k$-uniform) hypergraph to $\Phi^{\sigma}$. For variable $v\in V\setminus\Lambda$, let $\Phi_{v}^{\sigma}$ denote the connected component of $\Phi^{\sigma}$ to which $v$ belongs.
@@
When $S=\{v\}$ is a single vertex, we let $\mu_{v}=\mu_{\{v\}}$. Furthermore, given some partial assignment $\sigma\in\{0,1,\perp\}^{\Lambda}$ for $\Lambda\subseteq V$, if $S\cap\Lambda=\emptyset$, we let $\mu_{S}^{\sigma}(\cdot):=\mu_{S}(\cdot\mid\sigma)$ be the marginal distribution on $S$ conditioned on the partial assignment $\sigma$. When $S=V\setminus\Lambda$, we simply use $\mu^{\sigma}$ to denote $\mu^{\sigma}_{V\setminus\Lambda}$.